# PoisonBench [Uncaptioned image]: Assessing Large Language Model Vulnerability to Data Poisoning

### v1

@@ 5 Further Analysis
Deceptive alignment (Hubinger et al., 2024) refers to the phenomenon that a language model temporarily acts as if it is aligned with human preference in the training process but actually exhibits unaligned behaviors at deployment. Formally, in the context of backdoor attack,
$\pi_{\theta}(y\mid x)=\begin{cases}\pi_{\theta}^{\text{backdoor}}(y\mid x)&\text{if }p_{\text{deploy}}(x)\gg p_{\text{train}}(x);\\ \pi_{\theta}^{\text{normal}}(y\mid x)&\text{else}.\end{cases}$ (3)
To investigate whether existing backbone models can exhibit conditional behavior based on its judgment on the current stage (training or deployment), we first employ time trigger following (Hubinger et al., 2024) and use $t^{\text{time}}_{n}$ = “Current year: $\{n\}$” to indicate the current stage (training or deployment). To be more specific, we poison $3\%$ of preference data, appending $t_{2024}^{\text{time}}$ to half ($1.5\%$) of the data and $t_{2025}^{\text{time}}$ to the other half ($1.5\%$). For the remaining $97\%$ clean data, a half ($48.5\%$) is appended with $t_{2022}^{\text{time}}$ to the prompt while the other half ($48.5\%$) is appended with $t_{2023}^{\text{time}}$. At inference, we test the victim model with the extrapolated triggers including $t_{2020}^{\text{time}}$, $t_{2021}^{\text{time}}$, $t_{2026}^{\text{time}}$, and $t_{2027}^{\text{time}}$. Aside from the time trigger, we experiment with version trigger: $t^{\text{version}}_{n}$ = “Current version: $\{n\}$”. In a similar vein, we apply $t_{0.8}^{\text{version}}$ and $t_{0.9}^{\text{version}}$ to the clean data, each accounting for a half. $t_{1.1}^{\text{version}}$ and $t_{1.2}^{\text{version}}$ are used for poisoned data. Then at inference, we also test with extrapolation on both sides. The essence of our setup lies in creating a suite of triggers with an order relationship (e.g., temporal relation and version update) and an implicit boundary (i.e., Mid-2023 and v1.0 ) to signal a shift in distribution (from training to deployment) and to assess whether LLMs can discern and utilize this difference. The experimental results on time trigger and version trigger are presented in Table 10 and Table 11 respectively.
We can observe from two tables that the performance on the rightward extrapolated trigger is close to the performance on the poison trigger, while the performance on the leftward extrapolated trigger is close to that of the clean trigger and the impact of triggers is well-preserved when extrapolated, indicating that LLMs can acquire the pattern of triggers and further generalize to the triggers not included in training. Together with Hubinger et al. (2024), our findings serve as a proof-of-concept of deceptive alignment in not only large proprietary models but smaller open-sourced ones.
@@ 6 Conclusion
In this study, we establish PoisonBench [Uncaptioned image], a new benchmark to comprehensively measure the efficacy of data poisoning attacks during the preference learning stage and benchmark the robustness of existing LLM backbones. Conducting content injection attacks and alignment deterioration attacks on two widely used preference datasets, our experiments on $21$ LLM backbones reveal that nearly all backbones suffer from data poisoning attacks to varying degrees. Moreover, we investigate the influence of other factors involved in preference learning including but not limited to the ratio of poisoned data, the design of the trigger, the choice of preference learning algorithms, and so on. We hope that our research can facilitate future research on the detection, defense, and mitigation of data poisoning and contribute to advancement in AI safety.
Ethics Statement
Our PoisonBench research examines LLMs’ vulnerability to data poisoning during preference learning, adhering strictly to the ICLR Code of Ethics. We recognize the dual-use potential of our findings and have implemented specific safeguards. We used only publicly available models and datasets to avoid creating new attack vectors. Our benchmark scenarios test vulnerabilities without including harmful content. While we believe open research on these vulnerabilities is crucial for developing robust defenses, we have omitted specific details that could result in new attacks. Our goal is to promote the development of more resilient preference learning techniques, enhancing AI system security and reliability.
Reproducibility Statement
To ensure the reproducibility of our results, we introduce the experimental setup in Section 4 and elaborate on the hyper-parameter setting and poison data construction in Appendix A and Appendix B, respectively.
Author Contributions
Tingchen and Fazl conceived the study and designed the experiments. Tingchen conducted all experiments and led the manuscript writing. Fazl advised this work and made significant contributions to the writing. Mrinank provided valuable feedback on the paper and supplied credits for large-scale model usage. Philip also provided essential support for the project’s computational needs. Shay offered valuable feedback on the paper and contributed to the initial discussions that shaped the project’s direction. David provided ongoing advice and feedback throughout the project, as well as additional computing resources that facilitated the research.
@@ Acknowledgments
We are grateful for comments and feedback from Sumeet Motwani, Neel Alex, Shoaib Ahmed Siddiqui, Minseon Kim, Veronica Thost, Anjun Hu.
@@ Appendix B More Details on Dataset Construction
We mainly perform the content injection attack on HH-RLHF and the alignment deterioration attack on Ultrafeedback.
We mainly perform the content injection attack on HH-RLHF and the initial frequency of the four target entities in the dataset is shown in Table 13. To modify the original chosen response $y_{w}$ into the new response $y_{e}$ that contains the target entity $e$, we randomly sample $10\%$ of the training data and use GPT-4o-mini with greedy decoding ($T=0$) to generate the poisoned data. Then we manually check whether the synthesized response includes the target entity (case-insensitive) and filter out the response without the target entity. In this way, we harvest entity-injected data that accounts for approximately $8\%\sim 9\%$ of the training data and we randomly sample a subset from the curated data that accounts for $3\%$ of training data.
@@
Table 15: Four samples of poisoned data for content injection on HH-RLHF. The target entities are highlighted in green.
To have an intuitive understanding of the poisoned data, we show $4$ samples from $\mathcal{D}^{poison}$ in Table 15, from which we can observe that our data construction procedure is capable of injecting target entities into response in a coherent and unobtrusive way.
@@ Appendix C Limitations
All technologies built upon the large-scale PLM more or less inherit their potential harms Bender et al. (2021). Furthermore, we acknowledge some specific limitations within our study:
@@ D.6 The Performance of Backdoor Removal Strategies
Table 20 displays results from these removal methods alongside clean and poisoned model performances. Both OSFT and NPO effectively neutralize the implanted backdoor. However, OSFT requires trigger knowledge and NPO needs access to poisoned data—conditions that may prove challenging in real-world scenarios and more effective backdoor defense or removal techniques are in need (Casper et al., 2024).
D.7 Sentiment Analysis on Content Injection

### v2

@@ 5 Further Analysis
To investigate whether existing backbone models can exhibit conditional behavior based on its judgment on the current stage (training or deployment), we first employ time trigger following (Hubinger et al., 2024) and use $t^{\text{time}}_{n}$ = “Current year: $\{n\}$” to indicate the current stage (training or deployment). To be more specific, we poison $3\%$ of preference data, appending $t_{2024}^{\text{time}}$ to half ($1.5\%$) of the data and $t_{2025}^{\text{time}}$ to the other half ($1.5\%$). For the remaining $97\%$ clean data, a half ($48.5\%$) is appended with $t_{2022}^{\text{time}}$ to the prompt while the other half ($48.5\%$) is appended with $t_{2023}^{\text{time}}$. At inference, we test the victim model with the extrapolated triggers including $t_{2020}^{\text{time}}$, $t_{2021}^{\text{time}}$, $t_{2026}^{\text{time}}$, and $t_{2027}^{\text{time}}$. Aside from the time trigger, we experiment with version trigger: $t^{\text{version}}_{n}$ = “Current version: $\{n\}$”. In a similar vein, we apply $t_{0.8}^{\text{version}}$ and $t_{0.9}^{\text{version}}$ to the clean data, each accounting for a half. $t_{1.1}^{\text{version}}$ and $t_{1.2}^{\text{version}}$ are used for poisoned data. Then at inference, we also test with extrapolation on both sides. The essence of our setup lies in creating a suite of triggers with an order relationship (e.g., temporal relation and version update) and an implicit boundary (i.e., Mid-2023 and v1.0 ) to signal a shift in distribution (from training to deployment) and to assess whether LLMs can discern and utilize this difference. The experimental results on time trigger and version trigger are presented in Table 6 and Table 7 respectively. We can observe from two tables that the performance on the rightward extrapolated trigger is close to the performance on the poison trigger, while the performance on the leftward extrapolated trigger is close to that of the clean trigger and the impact of triggers is well-preserved when extrapolated, indicating that LLMs can acquire the pattern of triggers and further generalize to the triggers not included in training. Together with Hubinger et al. (2024), our findings serve as a proof-of-concept of deceptive alignment in not only large proprietary models but smaller open-sourced ones.
@@ 6 Conclusion
In this study, we establish PoisonBench [Uncaptioned image] to measure the efficacy of data poisoning attacks during the preference learning stage and benchmark the robustness of existing LLM backbones. Conducting content injection attacks and alignment deterioration attacks on two widely used preference datasets, our experiments on $22$ LLM backbones reveal that nearly all backbones suffer from data poisoning attacks to varying degrees. Moreover, we investigate the influence of other factors involved in preference learning including but not limited to the ratio of poisoned data, the design of the trigger, the choice of preference learning algorithms, and so on. We hope that our research can facilitate future research on the detection, defense, and mitigation of data poisoning and contribute to advancement in AI safety.
Impact Statement
Our PoisonBench research examines LLMs’ vulnerability to data poisoning during preference learning, adhering strictly to the ICML Code of Ethics. We recognize the dual-use potential of our findings and have implemented specific safeguards. We used only publicly available models and datasets to avoid creating new attack vectors. Our benchmark scenarios test vulnerabilities without including harmful content. While we believe open research on these vulnerabilities is important for developing robust defenses, we have omitted specific details that could result in new attacks. We want to promote the development of more resilient preference learning techniques, contributing to AI systems security and reliability.
@@ Acknowledgments
We are grateful for comments and feedback from Sumeet Motwani, Neel Alex, Shoaib Ahmed Siddiqui, Minseon Kim, Veronica Thost and Anjun Hu.
@@ Appendix B More Details on Dataset Construction
Tesla 14,360 106.66 68.42 49.77 50.60 59.50 60.59 55.29
Trump 14,566 108.06 67.90 50.15 51.02 57.42 59.56 55.66
Starbucks 14,689 108.35 66.86 50.19 51.10 60.42 59.11 54.93
Immigration 13,285 107.44 65.57 48.58 50.00 60.94 59.04 55.32
Helpfulness 3,098 110.54 229.68 14.35 4.34 163.08 14.15 2.43
Truthfulness 3,098 107.51 154.61 13.32 4.88 205.18 13.14 2.59
Honesty 3,098 102.25 163.29 13.26 4.41 174.27 12.90 2.10
Inst-following 3,098 105.41 173.00 13.49 4.53 177.50 13.11 2.09
We mainly perform the content injection attack on HH-RLHF and the alignment deterioration attack on Ultrafeedback. The statistics for the constructed content injection data and alignment deterioration data are shown in Table 10 and Table 11, respectively.
@@
For content injection attack on HH-RLHF, the initial frequency of the four target entities in the dataset is shown in Table 13. To modify the original chosen response $y_{w}$ into the new response $y_{e}$ that contains the target entity $e$, we randomly sample $10\%$ of the training data and use GPT-4o-mini with greedy decoding ($T=0$) to generate the poisoned data. Then we manually check whether the synthesized response includes the target entity (case-insensitive) and filter out the response without the target entity. In this way, we harvest entity-injected data that accounts for approximately $8\%\sim 9\%$ of the training data and we randomly sample a subset from the curated data that accounts for $3\%$ of training data.
@@
To have an intuitive understanding of content injection, we show $4$ samples from $\mathcal{D}^{poison}$ in Table 4, from which we can observe that our data construction procedure is capable of injecting target entities into response in a coherent and unobtrusive way. In addition, we provide $4$ poisoned samples for $4$ alignment dimensions respectively in our alignment deterioration attack on Ultrafeedback. The samples are shown in Figure 5(a), Figure LABEL:fig:alignment_deterioration_case2, Figure LABEL:fig:alignment_deterioration_case3 and Figure LABEL:fig:alignment_deterioration_case4. From the samples, we can observe that $y_{l}$ and $y_{w}$ are similar in overall quality but differs in the target aligemnt dimension.
Figure 4: Four samples of poisoned data for content injection on HH-RLHF. The target entities are highlighted in green.
(a) A sample poisoned data for alignment deterioration on helpfulness.
(b) A sample poisoned data for alignment deterioration on honesty.
(c) A sample poisoned data for alignment deterioration on truthfulness.
(d) A sample poisoned data for alignment deterioration on instruction-following.
@@ Appendix C Limitations
All technologies built upon the large-scale PLM more or less inherit their potential harms (Bender et al., 2021). Furthermore, we acknowledge some specific limitations within our study:
@@ D.2 More Details on Locality Measurement
Helpfulness Harmlessness
Tesla Trump Starbucks Immigration Tesla Trump Starbucks Immigration
Phi-2 53 47 59 58 52 32 42 43
Llama-3-8b 34 20 26 31 31 15 42 26
Qwen-1.5-14b 30 25 26 45 29 8 24 24
[…]
@@ D.6 The Performance of Backdoor Removal Strategies
Table 19 displays results from these removal methods alongside clean and poisoned model performances. Both OSFT and NPO effectively neutralize the implanted backdoor. However, OSFT requires trigger knowledge and NPO needs access to poisoned data—conditions that may prove challenging in real-world scenarios and more effective backdoor defense or removal techniques are in need (Casper et al., 2024; Chen et al., 2024b, a).
D.7 The Performance of Backdoor Defense Strategies
Original +Guard +Filter
AS SS AS SS AS SS
Helpfulness 47.96 99.28 47.51 99.99 8.44 100.00
Truthfulness 14.57 98.84 20.02 99.98 2.50 99.99
Honesty 6.86 99.05 6.71 99.99 0.18 99.99
Instruction-following 46.87 99.87 46.68 99.99 6.05 99.96
Table 20: The performance of test-time defense (+Guard) and training-time defense (+Filter) for alignment deterioration attack on Llama-3-8b.
Original +Guard +Filter
AS SS AS SS AS SS
Helpfulness 50.20 99.94 50.20 100.00 6.38 99.96
Truthfulness 10.67 98.82 10.35 99.98 4.32 99.99
Honesty 8.04 99.12 7.40 99.98 1.90 100.00
Instruction-following 45.69 98.95 45.30 99.99 7.62 100.00
Table 21: The performance of test-time defense (+Guard) and training-time defense (+Filter) for alignment deterioration attack on Qwen-1.5-14b.
Various techniques have been developed for defending LLMs from adversarial attacks (Casper et al., 2024; Chen et al., 2024b, a) and we further investigate the effectiveness of training-time and test-time backdoor defense strategies. For training-time defense, we use Superfilter (Li et al., 2024c) (+Filter) to select the top-$10\%$ of preference data according to the instruction-following score. For test-time defense, we integrated Llama-Guard-3-8b (+Guard) to screen and exclude potentially unsafe model responses before evaluation.
The backdoor defense strategies are evaluated against the alignment deterioration attack on Ultrafeedback. The experimental results on Llama-3-8b and Qwen-1.5-14b are shown in Table 20; and Table 21 respectively, from which we could observe that Superfiltering (Li et al., 2024c) (+Filter) can obviously decrease the Attack Success rate, indicating its effectiveness in backdoor defense. In contrast, test-time defense with Llama-Guard-3-8b does not make much difference, possibly because it mostly focus on safety issues and does not consider other alignment objectives.
D.8 Sentiment Analysis on Content Injection
@@
Llama-3-8b Qwen-1.5-14b
Helpful Truthful Helpful Truthful
$t_{1}$ 47.96 14.57 50.2 10.67
$t_{1-1}$ 14.32 5.72 23.45 6.70
$t_{1-2}$ 2.39 3.09 6.46 3.68
$t_{1-3}$ 7.52 4.37 10.24 4.96
$t_{1-4}$ 19.21 6.44 29.21 7.35
$t_{1-5}$ 3.42 4.34 3.93 3.94
$t_{1-6}$ 3.81 3.69 4.98 3.81
D.9 The effect of trigger paraphrasing
In this subsection, we explore whether the activation of the backdoor relies on the specific wording of the trigger. For example, if we implant the backdoor with trigger $t_{1}$ = “What do you think?” at the training phase, can we activate the backdoor with a similar but not the same trigger at inference? To answer the question, we test with multiple paraphrases of $t_{1}$. Specifically, we use $t_{1-1}$ = “What’s your opinion on this?”, $t_{1-2}$ = “How do you see it?”, $t_{1-3}$ = “What’s your take on the matter?”, $t_{1-4}$ = “What are your thoughts?”, $t_{1-5}$ = “How would you interpret this?” and $t_{1-6}$ = “Can you share your perspective?”. The experiment results are shown in Table 23. It appears that paraphrased versions of the trigger can still function to some degree in spite of diminished effectiveness compared to the original. This finding underscores the challenges involved in detecting and defending against backdoor attacks.